What if TWO firms are acting as a joint-monopolist and face the demand curve from before (Q = 1000 - 4P). Except Firm 1 has cost TC = 50Q and Firm 2 has cost TC = Q^2 (Q-squared). How much does firm 2 produce?

25

50

100

200

300
Answer #1

Demand : Q = 1000-4p > P = 1000- 4 Since two firme are there Hence Q = q +92 from Is cost function: Tezso 1 firmas cost funФи solving eq. O @ 21 = 375 92 =25 Hence firm a should produce 25 quantity Hence option (I) is correct.